OpenStudy (anonymous):

find d2y/dx2 in terms of x and y if y+1/x

OpenStudy (anonymous):

You're finding the second derivative, so to make your equation into y=x^(-1) to make it a lot easier. All you have to do is apply the power rule: y'=nx^(n-1) y=x^(-1) (so n would be -1) y'=-1x^(-2) to find the second derivative, just do the power rule again. (n=-2) y'=-1x^(-2) y"=2x^(-3) Hope that helps!